OpenStudy (anonymous):

Calculate the kinetic energy of a 5.0 kg ball traveling at 6 m/s

OpenStudy (anonymous):

\[KE = \frac{ 1 }{ 2 }mv^2\] You have m and you have v, just plug in and solve.

OpenStudy (anonymous):

so im plugging 5.0 times it by 2

OpenStudy (anonymous):

No. You're going to put 5 for m, and 6 for v, so what happens is: \[KE = \frac{ 1 }{ 2 }mv^2 = \frac{ 1 }{ 2 }*5*(6)^2 = \frac{ 1 }{ 2}* 5*36 = = \frac{ 1 }{ 2 }*180 = 90J\]

OpenStudy (anonymous):

oh okay i had part of that but thanks can you help me with another one please

OpenStudy (anonymous):

Sure, ask away? (BTW, click best response if this helped)

OpenStudy (anonymous):

k

OpenStudy (anonymous):

J of heat is required to raise the temperature of 100. grams of water by 10. degrees Celsius

OpenStudy (anonymous):

---- J of heat is required to raise the temperature of 100. grams of water by 10. degrees Celsius @Zelda

OpenStudy (anonymous):

@jmays14, no idea. I have no chem experience. I suggest asking this in the chem subtopic.

OpenStudy (anonymous):

k what about this question..... Calculate the density of an object with a volume of 5.0 cm3 and a mass of 13g

OpenStudy (anonymous):

@Zelda

OpenStudy (anonymous):

Density = m/v

OpenStudy (anonymous):

okay so how would i write it out as a problem to solve

OpenStudy (anonymous):

I just told you.

OpenStudy (anonymous):

Not much to it mate.

OpenStudy (anonymous):

@Zelda

OpenStudy (anonymous):

Yes?

OpenStudy (anonymous):

how would i write it out as a problem to solve

OpenStudy (anonymous):

there is NO problem to solve. DENSITY = M/V = MASS DIVIDED BY VOLUME = MASS/VOLUME. You just plug into the formula, and evaluate it.

OpenStudy (anonymous):

5.0/13

OpenStudy (anonymous):

@Zelda

OpenStudy (anonymous):

OMG. Honestly mate. Read it again, and again, and again. Even after I spelled it out. you still reversed it.

OpenStudy (anonymous):

i didnt revise it you said mass/ volume 13/5.0

OpenStudy (anonymous):

Yes, but you put 5/13... Anyhow, yes it is 13/5

OpenStudy (anonymous):

okay and i got 2.6 is that correct

OpenStudy (anonymous):

If that is what 13/5 is then yes. Sorry for being mad, got slightly frustrated there.

OpenStudy (anonymous):

Oh and your units are: g/cm^3

OpenStudy (anonymous):

huh now im lost on that one im gonna guess so it'll be 2.6^3= 7.8

OpenStudy (anonymous):

No, no no. Basically the unit for density is unit of mass/unit of volume So your \[D = 2.6 \frac{ g }{ cm^3 }\]

OpenStudy (anonymous):

oh ok do i have to really include that whole problem in my answer i cant just leave it as 2.6

OpenStudy (anonymous):

No, you always need units. Otherwise it could be 2.6 elephants!